OpenStudy (anonymous):

Factor completely: −2x^3 + 6x^2

OpenStudy (anonymous):

@welshfella

OpenStudy (freckles):

notice any common factor in -2x^3 and 6x^2

OpenStudy (freckles):

hint: \[-2x^3=(-1)(x)(2x^2) \\ 6x^2=(3)(2x^2)\]

OpenStudy (anonymous):

(2x^2)?

OpenStudy (freckles):

yes so factor that out

OpenStudy (anonymous):

-_- um.

OpenStudy (freckles):

what factors were leftover when we compared the factoziations of -2x^3 and 6x^2?

OpenStudy (anonymous):

6x^2=(3)?

OpenStudy (anonymous):

I have no idea, how to this ;-(

OpenStudy (freckles):

 you noticed we noticed the common factor is 2x^2?